# Harmonizing Generalization and Specialization: Uncertainty-Informed Collaborative Learning for Semi-supervised Medical Image Segmentation

### v1

@@ 4.6 Computational Efficiency and Resource Consumption
| Dim. | Method | Params(M) | MACs(G) | Memory(MB) | Latency(ms) |
| 2D | SAM [5] / MedSAM [6] | 93.74 | 371.46 | 2771.71 | 102.20 |
| UNet [31] Group | 1.81 | 3.01 | 145.96 | 1.86 | |
| nnU-Net [32] | 18.69 | 4.76 | 211.52 | 4.16 | |
| URPC [15] | 1.82 | 3.05 | 151.30 | 2.07 | |
| ABD [12] | 27.15 | 7.75 | 162.94 | 13.71 | |
| CPC-SAM [16] | 94.94 | 32.78 | 502.17 | 35.34 | |
| UnCoL (Ours) | 3.53 | 3.60 | 163.00 | 4.34 | |
| 3D | SAM-Med3D [7] | 100.51 | 186.44 | 1477.58 | 65.30 |
| VNet [33] Group | 9.44 | 97.91 | 810.74 | 14.93 | |
| nnU-Net [32] | 30.45 | 215.48 | 1134.37 | 35.68 | |
| Swin-UNETR [2] | 61.99 | 785.06 | 3980.14 | 194.09 | |
| MC-Net+ [22] | 15.25 | 298.96 | 1120.84 | 57.79 | |
| CauSSL [19] | 12.98 | 190.70 | 1576.71 | 43.13 | |
| Su et al. [14] | 12.35 | 198.83 | 1221.75 | 39.28 | |
| UnCoL (Ours) | 13.94 | 144.45 | 887.97 | 45.87 |
Table 7 compares the computational efficiency of different dimensional settings (2D and 3D) and learning paradigms. For fair comparison, methods sharing the same backbone are merged and reported as a group (e.g., MT/UAMT/BCP under U-Net or V-Net). Compared with lightweight baselines, UnCoL incurs only moderate computational overhead while achieving substantially higher accuracy (see Sec. 4.2), reflecting a favorable efficiency–accuracy trade-off. Moreover, it requires significantly less memory and latency than zero-shot or fine-tuned foundation models such as CPC-SAM, highlighting its practicality for efficient semi-supervised segmentation.

### v2

@@ 4.6 Computational Efficiency and Resource Consumption
| Dimension | Method | Params | MACs | Memory | Training Time | Inference Time |
| (M) | (G) | (MB) | (s/iteration) | (ms/sample) | ||
| 2D | SAM [18] / MedSAM [23] | - | 102.20 | |||
| UNet [30] Group | 0.16-0.51 | 1.86 | ||||
| nnU-Net [17] | 0.36 | 4.16 | ||||
| URPC [22] | 0.14 | 2.07 | ||||
| ABD [7] | 0.54 | 13.71 | ||||
| CPC-SAM [28] | 0.23 | 17.59 | ||||
| UnCoL (Ours) | 3.53 | 3.60 | 163.00 | 0.81 | 4.34 | |
| 3D | SAM-Med3D [34] | - | 65.30 | |||
| VNet [29] Group | 0.45-1.54 | 14.93 | ||||
| nnU-Net [17] | 1.13 | 35.68 | ||||
| Swin-UNETR [14] | 0.41 | 194.09 | ||||
| MC-Net+ [37] | 2.52 | 57.79 | ||||
| CauSSL [27] | 0.98 | 43.13 | ||||
| Su et al. [32] | 0.33 | 39.28 | ||||
| UnCoL (Ours) | 13.94 | 144.45 | 887.97 | 1.10 | 45.87 |
Table 7 compares the computational efficiency of different dimensional settings (2D and 3D) and learning paradigms, reporting parameter count, MACs, peak GPU memory, average training time per iteration (default batch size), and inference time per sample. Under the 2D setting, training time is measured on the OASIS dataset with 5% labeled data, and under the 3D setting on the Pancreas-CT dataset with 10% labeled data. For fair comparison, methods sharing identical backbones and comparable computational configurations are grouped together (e.g., MT/UAMT/BCP under the U-Net or V-Net backbone), and the reported training time reflects the range observed within each group. Compared with lightweight baselines, UnCoL incurs only moderate computational overhead while achieving substantially higher accuracy (see Sec. 4.2), demonstrating its practical effectiveness under constrained computational budgets. Moreover, it requires significantly less memory and inference time than zero-shot or fine-tuned foundation models such as CPC-SAM, highlighting its practicality for efficient semi-supervised segmentation.